The x-intercept of a line has a value of and a y-intercept of 7. Another line has a slope that is twice the slope of the first line and has a y-intercept that is 15 more than the first line’s y-intercept. What is the product of the coordinates in the solution to this system? Write a system of equations for this situation and find its solution.

Answer by ankor@dixie-net.com(22740)

Assume this means:
The x-intercept of a line has a value of 7 and a y-intercept of 7.
Find the slope (m)
7m + 7 = 0
7m = -7
m = -7/7
m = -1
then
y = -x + 7 is the first equation
:
Another line has a slope that is twice the slope of the first line and has a y-intercept that is 15 more than the first line’s y-intercept.
m = -2
y intercept: 15 + 7 = 22
y = -2x + 22 is the 2nd equation
:
solve the system by substitution y=y
-x + 7 = -2x + 22
+2x - x = 22 - 7
x = 15
find y
y = -15 + 7
y = -8
The solution to the system: 15,-8
:
What is the product of the coordinates in the solution to this system?
15 * -8 = -120
.